OpenStudy (anonymous):

Im the teachers favorite student and he said if i can solve this problem i can get 100 dollars......1804351 x 235-135425 x 10 to the fifth power + 1- 12438723562956918723847.

OpenStudy (anonymous):

-12438723563091920701361

OpenStudy (anonymous):

yours is wrong

OpenStudy (anonymous):

did you remember your order of operations

OpenStudy (anonymous):

dude its -1 is by itself do u no math

OpenStudy (anonymous):

Firs do the exponents which is 10 ^ 5

OpenStudy (anonymous):

10^5 = 10,000

OpenStudy (anonymous):

no thats not it

OpenStudy (anonymous):

I got that 10^5 = 100,000.

OpenStudy (anonymous):

no thats not it

OpenStudy (anonymous):

1804351 * 235 = 424022485 10^2 = 100000 135425 * 100000 = 13542500000 424022485 - 13542500000 = -13118477515 -13118477515 + 1 = -13118477516 -13118477516 - 12438723562956918723847 = -12438723562970037201363

OpenStudy (anonymous):

i got the answer

OpenStudy (anonymous):

I know it's not the final answer.

OpenStudy (anonymous):

i got the answer

OpenStudy (anonymous):

not u sue

OpenStudy (anonymous):

i wasnt talkin to u i was talking to garra

OpenStudy (anonymous):

oh so what's that answer

OpenStudy (anonymous):

you have to do it step by step using the PEMDAS

OpenStudy (anonymous):

look at my match chris...or you can post yours and i can show you what you did wrong

OpenStudy (anonymous):

ok i did it and my teacher said its right 1.2432981e+22

OpenStudy (anonymous):

my teacher said its right so thank u all but sorry

OpenStudy (anonymous):

That's not the real answer thought i mean u just put in the calculator right?

OpenStudy (anonymous):

I got the same thing cuz I just put the numbers in google lol

OpenStudy (anonymous):

but I was actually trying to do it by hand I guess i would've been wrong anyway :)

OpenStudy (anonymous):

No i worked it step by step i was working sense Saturday doing it

OpenStudy (anonymous):

show me your math chris, i'll show you where u went wrong xD look at my answer

OpenStudy (anonymous):

well ur right sue

OpenStudy (anonymous):

good job sue

OpenStudy (anonymous):

i dont need ur answer gaara i got the money

OpenStudy (anonymous):

Well @gaara438125 looks correct

OpenStudy (anonymous):

well yall can stop typing now